If Ann works 5 1/2 hrs., Mary works 6 1/3 hrs., and John works 4 1/4 hrs. How many combined hours have they worked?

Answer by rfer(16322)

5 1/2+6 1/3+4 1/4=
5 6/12+6 4/12+4 3/12=15 13/12= 16 1/12